OpenStudy (anonymous):

The length of the base of an isosceles triangle is 30 m. The angle opposite to the base is 32 degrees. Find the perimeter of the triangle. Help?

OpenStudy (anonymous):

Draw a triangle ABC as follows:- Base is BC = 30 A = 32° B = C = 74° Sine Rule 30 / sin 32° = AC / sin 74° AC = 30 sin 74° / sin 32° AC = 54.4 AB = 54.4 P = 108.8 + 30 m P = 138.8 m <3Hope it helpssss

OpenStudy (anonymous):

how did you get 74? :$

OpenStudy (anonymous):

Drop a perpendicular from the 32 degree angle to the base. Now you have 2 triangles, each of which has angles of 16 degrees, 90 degrees and (by difference from 180) 74 degrees. Cos 74 = .2756 = 15/x where x is one of the 2 equal legs of the isosceles triangle. x = 54.427 Perimeter = 2 X 54.427 + 30 = 138.853

OpenStudy (anonymous):

ohh why did you put 15/x? Other than that thank you so much! :)

OpenStudy (anonymous):

x is one of the 2 equal legs of the isosceles triangle. yourwel! :D

OpenStudy (anonymous):

Consider the height as h, two equal sides as b and the base as a. Then use this formula - h=sqrt(b^2-1/4a^2).

OpenStudy (anonymous):

h = b*cos(1/2 of 32) Let x be half of base (a) ie 30 = x = b*sin(1/2 of 32), i.e h=b*cos16 and 30=b*sin16 Solve for b Now perimeter is sum of all sides i.e a+2b i.e 30+2b